# U-PINet: Physics-Informed Hierarchical Learning for Accurate and Fast 3D RCS Prediction

### v2

@@ II-B Methodological Foundation Inspired by MLFMA
From a structural viewpoint, MLFMA is essentially built on the interplay between near-field accurate interaction, far-field multilevel evaluation, and the residual-driven iterative solver of the EFIE, forming a unified computational flow that preserves full electromagnetic physics while reducing complexity.
First, pairs of basis functions whose supports are geometrically close are collected in the near-neighbor list $\mathcal{N}$, and its contributions are computed exactly to form the near-field impedance block $\mathbf{Z}_{\text{near}}$. This block contributes the term $\mathbf{Z}_{\text{near}}\boldsymbol{I}$, which captures local self and mutual coupling in a purely MoM.
Second, the far-field interactions are organized on an $L$-level octree [38]. At each level $l$, the outgoing radiation patterns $\boldsymbol{\beta}^{(l)}$ of source groups are first aggregated to the parent level through a phase-shifted combination of child patterns. For each group, well-separated interactions are handled by a translation operator $\mathcal{T}_{l}$, which maps the aggregated patterns $\boldsymbol{\beta}^{(l)}$ of groups in the interaction list $\mathcal{I}$ to incoming fields $\boldsymbol{\alpha}^{(l)}_{\mathrm{trans}}$.
At the end of this multilevel evaluation, the finest-level translated field $\boldsymbol{\alpha}^{(L)}$ yields the far-field contribution of the MoM matrix–vector product, denoted as $\mathbf{Z}_{\mathrm{far}}\boldsymbol{I}$. Together with the exactly computed near-field term $\mathbf{Z}_{\mathrm{near}}\boldsymbol{I}$, MLFMA reconstructs the full matrix–vector multiplication
@@
Third, this residual is then used by an iterative Krylov solver (e.g., GMRES [39] or BiCGStab [40]) to update the current vector $\boldsymbol{I}$ until the boundary condition is satisfied, completing the residual-driven enforcement of the EFIE within the MLFMA framework. After convergence, the solved coefficient vector $\boldsymbol{I}$ is substituted back into the RWG expansion to recover the induced surface current $\boldsymbol{J}(\cdot)$. The scattered field is then evaluated via (3), and the desired bistatic RCS signature $\sigma(\theta,\phi)$ is obtained by (4).
Algorithm 1 summarizes the physically structured data flow of MLFMA, which serves as a reference model for the design of our physics-guided near–far learning architecture in U-PINet.
II-C Physics-Informed Design Mechanism of U-PINet
To enable efficient and physically consistent RCS evaluation, U-PINet follows the near–far organization of MLFMA while directly targeting accurate far-field scattering responses, as illustrated in Fig. 2. The resulting near- and far-field components are fused into the predicted current $\boldsymbol{J}_{\mathrm{pred}}$, from which the scattered field and RCS signature $\sigma(\theta,\phi)$ are obtained; a boundary loss constructed from the dyadic Green’s function is used to impose the EFIE constraint during training. Algorithm 2 summarizes this end-to-end near–far scheme.
II-C1 Near-Field EM Interactions Modeling
To model near-field interactions in a MoM-consistent manner, we construct a graph on the discretized surface points. For each observation point $\boldsymbol{p}_{i}\in\mathcal{P}$, its neighborhood is defined as:
@@ II-C1 Near-Field EM Interactions Modeling
where $r_{\mathrm{nf}}$ denotes the near-field interaction radius, which is chosen proportional to the local electrical size to capture strong short-range electromagnetic coupling.
collects geometrically close source points, analogous to the near-neighbor list used in the construction of $\mathbf{Z}_{\mathrm{near}}$. On each edge $(i,j)$, we define a physics-guided edge feature:s
@@
where $\mathcal{G}(\cdot)$ encodes the dependence of the Green’s function on inter-point distance and normal alignment. This design turns the edge attributes into a learnable surrogate of the self and mutual impedance terms in MoM, while preserving its geometric and electromagnetic structure.
On the constructed graph, the near-field module of U-PINet updates the current at node $\boldsymbol{p}_{i}$ through an iterative message-passing scheme
@@
where $h_{t}(\cdot)$ is a learnable local kernel at iteration $t$ that transforms the message from neighbor $j$ to node $i$ based on the edge feature $\boldsymbol{w}_{ij}$ and the neighbor current $\boldsymbol{J}^{(t)}(\boldsymbol{p}_{j})$. Structurally, (8) follows the same aggregation pattern as the MoM near-field operator
@@
where the contribution at index $i$ is obtained by summing the weighted influence of its near-field neighbors.
If we restrict $h_{t}$ to be linear in $\boldsymbol{J}^{(t)}(\boldsymbol{p}_{j})$, the message term can be written as
@@
with $K_{t}(i,j)$ acting as an effective, data-driven near-field coefficient determined by the physics-guided edge feature $\boldsymbol{w}_{ij}$.
In this view, the update in (8) is precisely a graph neural network (GNN) [41] message-passing block. At each iteration $t$, the node associated with $\boldsymbol{p}_{i}$ aggregates the current estimates $\{\boldsymbol{J}^{(t)}(\boldsymbol{p}_{j})\}_{j\in\mathcal{N}(i)}$ of its near-field neighbors, weighs each contribution through the physics-guided kernel $h_{t}(w_{ij},\cdot)$, and accumulates them to form a local correction. Structurally, this aggregation mirrors the MoM near-field matrix–vector product, where the $i$-th row of $\mathbf{Z}_{\mathrm{near}}$ collects and sums the influence of all neighboring basis functions on the $i$-th degree of freedom.
As a result, the near-field module can be viewed as a MoM-informed GNN operator that performs physics-guided message passing on the mesh graph. It remains computationally efficient and adaptable across different geometries and tessellations, while staying tightly anchored to the underlying EM physics through the MoM-consistent neighborhood $\mathcal{N}(i)$ and the Green’s-function-based edge features $\{w_{ij}\}$.
II-C2 Far-Field EM Interactions Modeling
To highlight how U-PINet is informed by the MLFMA structure, we first rewrite the far-field stages of Algorithm 1 in a compact operator form. On level $l$ of the octree, the aggregation, translation, and disaggregation steps can be summarized as
@@ II-C2 Far-Field EM Interactions Modeling
where $\mathcal{A}_{l}(\cdot)$ denotes the level-$l$ aggregation operator that interpolates and phase-shifts child radiation patterns into the parent pattern, $\mathcal{T}_{l}(\cdot)$ is the level-dependent translation operator that maps outgoing patterns $\boldsymbol{\beta}^{(l)}$ of well-separated groups to incoming fields $\boldsymbol{\alpha}^{(l)}_{\mathrm{trans}}$, and $\mathcal{D}_{l}(\cdot,\cdot)$ performs disaggregation (anterpolation and phase compensation) from coarse to fine levels. At the finest level, the accumulated incoming field is finally projected back to the testing basis functions to yield the far-field contribution
@@
with $\mathcal{R}(\cdot)$ denoting the radiation operator that recovers the far-field part of the MoM matrix–vector product.
@@
To mimic the multilevel far-field translation of MLFMA in a learnable manner, U-PINet introduces a cross-scale attention block [42], at each hierarchy level. Given the level-$k$ features $\boldsymbol{S}^{(k)}$ and the coarser representation $\tilde{\boldsymbol{S}}^{(k-1)}$, the global translation step is written as
@@
where $\boldsymbol{\Phi}_{k}(\cdot,\cdot)$ denotes the learnable cross-scale translation block and $\mathcal{U}_{k}(\cdot)$ is an unpooling operator that lifts the coarse-level representation $\tilde{\boldsymbol{S}}^{(k-1)}$ to level $k$. In this formulation, $\boldsymbol{S}^{(k)}$ plays the role of level-$k$ observation groups, while $\mathcal{U}_{k}(\tilde{\boldsymbol{S}}^{(k-1)})$ provides a global source context analogous to the multigroup far-field patterns in MLFMA.
Concretely, the block $\boldsymbol{\Phi}_{k}$ is implemented as a cross-attention layer. Let
@@
where $\boldsymbol{Q}^{(k)}$, $\boldsymbol{K}^{(k)}$, and $\boldsymbol{V}^{(k)}$ denote the query, key, and value feature embeddings at level $k$, respectively, and $\mathbf{W}_{Q}^{(k)}$, $\mathbf{W}_{K}^{(k)}$, and $\mathbf{W}_{V}^{(k)}$ are trainable linear projection matrices that map the hierarchical source representations into a shared attention subspace for modeling cross-scale electromagnetic interactions. For an observation group $i$ and a source group $m$, the attention weight is given by
$a^{(k)}_{i,m}=\mathrm{softmax}_{m}\!\big(\langle\boldsymbol{Q}^{(k)}_{i},\boldsymbol{K}^{(k)}_{m}\rangle\big)=\frac{\exp\big(\langle\boldsymbol{Q}^{(k)}_{i},\boldsymbol{K}^{(k)}_{m}\rangle\big)}{\sum_{m^{\prime}}\exp\big(\langle\boldsymbol{Q}^{(k)}_{i},\boldsymbol{K}^{(k)}_{m^{\prime}}\rangle\big)}$, (20)
where $\langle\cdot,\cdot\rangle$ denotes the standard inner product between two feature vectors, $\exp(\cdot)$ is the exponential function, and $\mathrm{softmax}_{m}(\cdot)$ denotes the softmax normalization over all source groups indexed by $m$ [42]. The updated representation of group $i$ is obtained as
@@
In this view, the attention weights $a^{(k)}_{i,m}$ function as a learned analogue of the MLFMA translation kernel $T_{l}(i,m)$, while the value vectors $\boldsymbol{V}^{(k)}_{m}$ serve as learnable far-field source patterns; together with the hierarchy, the block $\boldsymbol{\Phi}_{k}$ provides a physics-guided neural counterpart of the classical multilevel far-field translation process.
Figure 3: Workflow of the proposed U-PINet for physics-informed surface current prediction.
The far-field current contribution is then recovered through a hierarchical disaggregation step,
@@
Comparing (11)-(14) with (15)-(22), U-PINet preserves the same near–far information flow as MLFMA: fine-to-coarse aggregation, level-wise translation, and coarse-to-fine reconstruction. The crucial change is that the fixed analytical operators $(\mathcal{A}_{l},\mathcal{T}_{l},\mathcal{D}_{l},\mathcal{R})$ are replaced by learnable counterparts $(\boldsymbol{\Gamma}_{k},\boldsymbol{\Phi}_{k},\boldsymbol{\Psi}_{k})$.
II-C3 Physics Residual Error
@@
The residual enforcement in U-PINet mirrors the residual-driven update mechanism found in the classical MLFMA pipeline. Specifically, the network outputs a predicted surface current density, decomposed into near- and far-field components as $\boldsymbol{J}_{\mathrm{pred}}=\boldsymbol{J}_{\mathrm{near}}+\boldsymbol{J}_{\mathrm{far}}$. Using this prediction, the scattered electric field is reconstructed via the dyadic Green’s function integral defined in (3). To ensure physical consistency, the PEC boundary condition formulated in (1) is directly enforced by minimizing the physics-informed residual loss:
@@
$\mathcal{L}_{\mathrm{phys}}=\mathbb{E}_{\boldsymbol{p}_{i}\in\mathcal{P}}\Big[\big\|\hat{\boldsymbol{n}}_{i}\times\big(\boldsymbol{E}^{\mathrm{inc}}(\boldsymbol{p}_{i})+\boldsymbol{E}^{\mathrm{sca}}(\boldsymbol{p}_{i};\boldsymbol{J}_{\mathrm{pred}})\big)\big\|^{2}\Big],$ (24)
@@
where $\mathcal{P}$ denotes the set of collocation points on the discretized surface.
@@
In contrast to the classical MLFMA, where the residual drives a Krylov subspace solver (e.g., GMRES or BiCGStab) to iteratively update the current coefficients, U-PINet utilizes $\mathcal{L}_{\mathrm{phys}}$ to optimize the trainable parameters of the hierarchical near–far architecture. This training process ensures that a single forward pass produces a current distribution $\boldsymbol{J}_{\mathrm{pred}}$ that approximately satisfies the EFIE. Consequently, the physics loss in U-PINet plays a role analogous to the residual vector in solver-based methods, effectively closing the loop between near-field graph modeling, multilevel far-field coupling, and electromagnetic boundary-condition enforcement.
@@
Once the network is trained, the resulting $\boldsymbol{J}_{\mathrm{pred}}$ is employed to evaluate the far-field scattering response via (3), and the final bistatic RCS signature $\sigma(\theta,\phi)$ is computed according to the far-field definition in (4).
III The Proposed U-PINet Framework
In this section, we present the overall architecture of the proposed U-PINet framework for physics-informed electromagnetic scattering modeling. As shown in Figure 3, the framework mirrors the classical MLFMA procedure in a multi-stage, end-to-end learnable manner. The pipeline takes the object geometry, surface normals, and incident field as inputs and outputs the predicted surface current.
III-A Near-Field Graph Encoder Module

### v4

@@ II-B The MLFMA Framework
From a structural viewpoint, MLFMA is essentially built on the interplay between near-field accurate interaction, far-field multilevel evaluation, and the residual-driven iterative solver of the EFIE, forming a unified computational flow that preserves full EM physics while reducing complexity. First, pairs of basis functions whose supports are geometrically close are collected in the near-neighbor list $\mathcal{N}$, and its contributions are computed exactly to form the near-field impedance block $\mathbf{Z}_{\text{near}}$. This block contributes the term $\mathbf{Z}_{\text{near}}\boldsymbol{I}$, which captures local self and mutual coupling in a purely MoM. Second, the far-field interactions are organized on an $L$-level octree [39]. At each level $l$ ($l=1,\dots,L$), the outgoing radiation patterns $\boldsymbol{\beta}^{(l)}$ of source groups are first aggregated to the parent level through a phase-shifted combination of child patterns. For each group, well-separated interactions are handled by a translation operator $\mathcal{T}_{l}$, which maps the aggregated patterns $\boldsymbol{\beta}^{(l)}$ of groups in the interaction list $\mathcal{I}$ to incoming fields $\boldsymbol{\alpha}^{(l)}_{\mathrm{trans}}$. At the end of this multilevel evaluation, the finest-level translated field $\boldsymbol{\alpha}^{(L)}$ yields the far-field contribution of the MoM matrix–vector product, denoted as $\mathbf{Z}_{\mathrm{far}}\boldsymbol{I}$. Together with the exactly computed near-field term $\mathbf{Z}_{\mathrm{near}}\boldsymbol{I}$, MLFMA reconstructs the full matrix–vector multiplication
@@
Finally, this residual is then used by an iterative Krylov solver (e.g., GMRES [40] or BiCGStab [41]) to update the current vector $\boldsymbol{I}$ until the boundary condition is satisfied, completing the residual-driven enforcement of the EFIE within the MLFMA framework. After convergence, the solved coefficient vector $\boldsymbol{I}$ is substituted back into the RWG expansion to recover the induced surface current $\boldsymbol{J}(\cdot)$. The scattered field is then evaluated via (3), and the desired bistatic RCS signature $\sigma(\theta,\phi)$ is obtained by evaluating expression (4).
Algorithm 1 summarizes MLFMA’s physically structured data flow, which serves as a reference model for the design of our physics-guided near–far learning architecture in U-PINet.
III The Proposed U-PINet Framework
To enable efficient and physically consistent RCS evaluation, the proposed U-PINet, whose overall architecture is presented in the following Section IV, capitalizes on the near–far organization of MLFMA while directly targeting accurate far-field scattering responses, as illustrated in Fig. 2. Algorithm 2 summarizes the overall procedure, and the following subsections describe its key steps: Near-Field Interaction Learning corresponds to Steps 1 and 2; Far-Field Interaction Learning corresponds to Stages 3–5; Physics-Informed Residual Error corresponds to Step 6; and the final scattered-field/RCS evaluation corresponds to Step 7.
III-1 Near-Field Interaction Learning
To model near-field interactions in a MoM-consistent manner, we construct a graph on the discretized surface points. For each surface node (testing point) $\boldsymbol{p}_{i}\in\mathcal{P}$, its neighborhood is defined as follows:
@@ III-1 Near-Field Interaction Learning
where $r_{\mathrm{nf}}$ denotes the near-field interaction radius chosen proportional to the local electrical size to capture strong short-range EM coupling. This expression implies that $\mathcal{N}(i)$ gathers the geometrically nearby points that dominate self/mutual coupling, mirroring the near-field interaction list in MoM-based discretizations. On each edge $(i,j)$, we define a physics-guided edge feature as follows:
@@
where $\mathcal{G}(\cdot)$ encodes the dependence of the Green’s function on inter-point distance and normal alignment. This design turns the edge attributes into a learnable surrogate of the self and mutual impedance terms in MoM, while preserving its geometric and EM structure.
On the constructed graph, the near-field module of U-PINet updates the current at node $\boldsymbol{p}_{i}$ through an iterative message-passing scheme as:
@@
where $h_{t}(\cdot)$ is a learnable local kernel at iteration $t$ that transforms the message from neighbor $j$ to node $i$ based on the edge feature $\boldsymbol{w}_{ij}$ and the neighbor current $\boldsymbol{J}^{(t)}(\boldsymbol{p}_{j})$. Structurally, (8) follows the same aggregation pattern as the MoM near-field operator, i.e.:
@@
where the contribution at index $i$ is obtained by summing the weighted influence of its near-field neighbors. It is noted that, if we restrict $h_{t}$ to be linear in $\boldsymbol{J}^{(t)}(\boldsymbol{p}_{j})$, the message term can be written as follows:
@@
where $K_{t}(i,j)$ acts as an effective, data-driven near-field coefficient determined by the physics-guided edge feature $\boldsymbol{w}_{ij}$.
In this view, the update in (8) is precisely a graph neural network (GNN) [42] message-passing block. At each iteration $t$, the node associated with $\boldsymbol{p}_{i}$ aggregates the current estimates $\{\boldsymbol{J}^{(t)}(\boldsymbol{p}_{j})\}_{j\in\mathcal{N}(i)}$ of its near-field neighbors, weighs each contribution through the physics-guided kernel $h_{t}(\boldsymbol{w}_{ij},\cdot)$, and accumulates them to form a local correction. Structurally, this aggregation mirrors the MoM near-field matrix–vector product, where the $i$-th row of $\mathbf{Z}_{\mathrm{near}}$ collects and sums the influence of all neighboring basis functions on the $i$-th degree of freedom. As a result, the near-field module can be viewed as a MoM-informed GNN operator that performs physics-guided message passing on the mesh graph. It remains computationally efficient and adaptable across different geometries and tessellations, while staying tightly anchored to the underlying EM physics through the MoM-consistent neighborhood $\mathcal{N}(i)$ and the Green’s-function-based edge features $\boldsymbol{w}_{ij}’s$.
III-2 Far-Field Interaction Learning
To highlight how U-PINet is informed by the MLFMA structure, we first rewrite the far-field stages of Algorithm 1 in a compact operator form. On level $l$ of the octree, the aggregation, translation, and disaggregation steps can be summarized as follows:
@@ III-2 Far-Field Interaction Learning
where $\mathcal{A}_{l}(\cdot)$ denotes the level-$l$ aggregation operator that interpolates and phase-shifts child radiation patterns into the parent pattern, $\mathcal{T}_{l}(\cdot)$ is the level-dependent translation operator that maps outgoing patterns $\boldsymbol{\beta}^{(l)}$ of well-separated groups to incoming fields $\boldsymbol{\alpha}^{(l)}_{\mathrm{trans}}$, and $\mathcal{D}_{l}(\cdot,\cdot)$ performs disaggregation (anterpolation and phase compensation) from coarse to fine levels. At the finest level, the accumulated incoming field is finally projected back to the testing basis functions to yield the far-field contribution:
@@
where $\mathcal{R}(\cdot)$ represents the radiation operator that recovers the far-field part of the MoM matrix–vector product.
@@
To mimic the multilevel far-field translation of MLFMA in a learnable manner, U-PINet introduces a cross-scale attention block [43] at each hierarchy level. Given the level-$k$ features $\boldsymbol{S}^{(k)}$ and the coarser representation $\tilde{\boldsymbol{S}}^{(k-1)}$, the global translation step is written as follows:
@@
where $\boldsymbol{\Phi}_{k}(\cdot,\cdot)$ denotes the learnable cross-scale translation block and $\mathcal{U}_{k}(\cdot)$ is an unpooling operator that lifts the coarse-level representation $\tilde{\boldsymbol{S}}^{(k-1)}$ to level $k$. In this formulation, $\boldsymbol{S}^{(k)}$ plays the role of level-$k$ testing groups, while $\mathcal{U}_{k}(\tilde{\boldsymbol{S}}^{(k-1)})$ provides a global source context analogous to the multigroup far-field patterns in the MLFMA framework. Concretely, the block $\boldsymbol{\Phi}_{k}$ is implemented as a cross-attention layer. Let
@@
where $\boldsymbol{Q}^{(k)}$, $\boldsymbol{K}^{(k)}$, and $\boldsymbol{V}^{(k)}$ denote the query, key, and value feature embeddings at level $k$, respectively, and $\mathbf{W}_{Q}^{(k)}$, $\mathbf{W}_{K}^{(k)}$, and $\mathbf{W}_{V}^{(k)}$ are trainable linear projection matrices that map the hierarchical source representations into a shared attention subspace for learning cross-scale EM interactions. For an observation group $i$ and a source group $m$, the attention weight is given by the following expression:
$a^{(k)}_{i,m}\triangleq\mathrm{softmax}_{m}\!\big(\langle\boldsymbol{Q}^{(k)}_{i},\boldsymbol{K}^{(k)}_{m}\rangle\big)=\frac{\exp\big(\langle\boldsymbol{Q}^{(k)}_{i},\boldsymbol{K}^{(k)}_{m}\rangle\big)}{\sum_{m^{\prime}}\exp\big(\langle\boldsymbol{Q}^{(k)}_{i},\boldsymbol{K}^{(k)}_{m^{\prime}}\rangle\big)}$, (20)
where $\langle\cdot,\cdot\rangle$ denotes the standard inner product between two feature vectors, $\exp(\cdot)$ is the exponential function, and $\mathrm{softmax}_{m}(\cdot)$ denotes the softmax normalization over all source groups indexed by $m$ [43]. The updated representation of group $i$ is then obtained as follows:
@@
Figure 3: The architecture of the proposed U-PINet for physics-informed surface current prediction.
In this way, the attention weights $a^{(k)}_{i,m}$ function as a learned analogue of the MLFMA translation kernel $T_{l}(i,m)$, while the value vectors $\boldsymbol{V}^{(k)}_{m}$ serve as learnable far-field source patterns; together with the hierarchy, the block $\boldsymbol{\Phi}_{k}$ provides a physics-guided neural counterpart of the classical multilevel far-field translation process. The far-field current contribution is then recovered through a hierarchical disaggregation step,
@@
Comparing (11)-(14) with (15)-(22), U-PINet preserves the same near–far information flow as MLFMA: fine-to-coarse aggregation, level-wise translation, and coarse-to-fine reconstruction. The crucial change offered by U-PINet is that the fixed analytical operators $(\mathcal{A}_{l},\mathcal{T}_{l},\mathcal{D}_{l},\mathcal{R})$ are now replaced by learnable counterparts $(\boldsymbol{\Gamma}_{k},\boldsymbol{\Phi}_{k},\boldsymbol{\Psi}_{k})$.
III-3 Physics-Informed Residual Error
@@
The residual enforcement in U-PINet mirrors the residual-driven update mechanism found in the classical MLFMA pipeline. Specifically, the proposed network outputs a predicted surface current density, decomposed into near- and far-field components as $\boldsymbol{J}_{\mathrm{pred}}=\boldsymbol{J}_{\mathrm{near}}+\boldsymbol{J}_{\mathrm{far}}$. Using this prediction, the scattered electric field is then reconstructed via the dyadic Green’s function integral as expressed in (3). To ensure physical consistency, the PEC boundary condition formulated in (1) is directly enforced by minimizing the following physics-informed residual loss:
@@
$\mathcal{L}_{\mathrm{phys}}=\mathbb{E}_{\mathcal{P}}\left[\left\|\hat{\boldsymbol{n}}_{i}\times\left(\boldsymbol{E}^{\mathrm{inc}}(\boldsymbol{p}_{i})+\boldsymbol{E}^{\mathrm{sca}}(\boldsymbol{p}_{i};\boldsymbol{J}_{\mathrm{pred}})\right)\right\|^{2}\right],$ (24)
@@
where $\mathcal{P}$ denotes the set of collocation points on the discretized surface, i.e., $\boldsymbol{p}_{i}\in\mathcal{P}$ $\forall i$, over which the expactation is taken.
@@
In contrast to the classical MLFMA, where the residual drives a Krylov subspace solver (e.g., GMRES or BiCGStab) to iteratively update the current coefficients, U-PINet utilizes $\mathcal{L}_{\mathrm{phys}}$ to optimize the trainable parameters of the hierarchical near–far architecture. This training process ensures that a single forward pass produces a current distribution $\boldsymbol{J}_{\mathrm{pred}}$ that approximately satisfies the EFIE. Consequently, the physics-informed loss in U-PINet plays a role analogous to the residual vector in methods based on iterative solvers, effectively closing the loop between near-field graph modeling, multilevel far-field coupling, and EM boundary-condition enforcement.
@@
Finally, once the network is trained, the resulting $\boldsymbol{J}_{\mathrm{pred}}$ is employed to evaluate the far-field scattering response via (3), and the final bistatic RCS signature $\sigma(\theta,\phi)$ is computed according to the far-field definition in (4).
IV The U-PINet Neural Network Architecture
In this section, we present the architecture of the proposed artificial neural network for physics-informed EM scattering reconstruction, where the induced surface current constitutes the primary learning target. As shown in Fig. 3, U-PINet takes the object geometry, surface normals, and incident field as inputs, and then outputs the predicted surface current $\boldsymbol{J}_{\mathrm{pred}}$ in an end-to-end, multi-stage manner that mirrors the near–far organization of MLFMA. The subsequent physics-based mapping from $\boldsymbol{J}_{\mathrm{pred}}$ to the scattered field and the RCS signature, together with the physics-residual supervision used during training, is illustrated separately in Fig. 5.
IV-A Near-Field Graph Encoder Module